OpenStudy (anonymous):

What volume of 0.0962 M NaOH is required to exactly neutralize 10.00 ml of 0.128 M HCl?

OpenStudy (callisto):

no. of mole of HCl = 0.01 x 0.128 = 1.28x10^3 no of mole of NaOH = no of mole of HCl =1.28x10^3 vol of NaOH required = 1.28x10^3/0.0962 x1000 = 13.3ml